OpenStudy (anonymous):

A baseball pitcher throws a ball with a speed of 41 m/s. Estimate the average acceleration of the ball during the throwing motion. In throwing the baseball, the pitcher accelerates the ball through a displacement of about 3.5 m, from behind the body to the point where it is released.

OpenStudy (anonymous):

v= u+at u= 0 ,v= 41m/s s= 1/2 a t^2 t= sqrt(2s/a) given s (distance)= 3.5m 41= a*t 41= a* sqrt(2s/a) 41= sqrt(2as) 1681 = 2*3.5 *a a= 24.14m/s^2

OpenStudy (anonymous):

a= 240.14m/s^2

OpenStudy (anonymous):

The concepts that are can be used in this problem are: v = v0 + at x = x0 + v0t + 1/2(at^2) v^2 = V0^2 + 2a(x - x0) v - velocity a-constant acceleration t - time x - distance

OpenStudy (anonymous):

How did you determine which concept to use?

OpenStudy (anonymous):

Use the 3rd equation. V(final) is 41m/s, v(initial) is 0m/s because the baseball was accelerated from rest. The quantity (x-x0) is the displacement, 3.5m. Solving for a yields: a=(Vf^2 - Vi^2)/(2(x-x0)=(41^2-0)/(2*3.5)=240m/s^2 :)